OpenStudy (anonymous):

If you have 400.0 mL of water at 25.00 °C and add 140.0 mL of water at 95.00 °C, what is the final temperature of the mixture? Use 1.00 g/mL as the density of water.

OpenStudy (anonymous):

\[(mass)(\Delta T)(C_s)=(C_s)(\Delta T)(mass)\]

OpenStudy (anonymous):

\[(140.0g)(95.00-x)(4.184)=(4.184)(x-25.00)(400.0g)\] Solve for x

OpenStudy (anonymous):

55647.2-585.76x=x1673.6x-41840 97487.2=2259.36x x=43.15

OpenStudy (anonymous):

where did you get the 4.184?

OpenStudy (anonymous):

it's the specific heat capcity for water.

OpenStudy (anonymous):

ah thanks